# Inflationary attractors and radiative corrections in light of ACT

### v1

@@ III.2 Predictions for $(n_{s},r)$
Figure 2: Comparison of $n_{s}$ and $r$ predictions for $\xi$-attractors (left) and $\alpha$-attractor T and E-models (right). Here we see that percent level quantum corrections can move the $\xi$-attractors off of their strong coupling attractor and comfortably within the latest ACT posteriors for the central $N=55$ number of e-foldings. Additionally, we see that sub-percent level quantum corrections can induce fairly substantial shifts in the predictions for $\alpha$-attractors at small $\alpha$, where we have included the predictions spanning $N\in[50,60]$ for both T and E-models.
@@
As shown by [45, 28] for Higgs-like inflation in both the Palatini and metric variations, as $\delta$ is increased, the model begins to approach a new attractor of linear inflation $V(\varphi)\propto\varphi$. The same holds for the rest of the $\xi$-attractors considered here (in the metric variation). Choosing $\delta=0.2$ and $\log_{10}\xi\in[-8,5]$, one can see in Fig. 3 that they all begin to approach the predictions for linear inflation. As before, the higher polynomials approach the attractor more slowly in both $\delta$ and $\xi$ as their potentials are more robust against quantum corrections.
Figure 3: As the loop correction term $\delta$ is made larger, the entire family of $\xi$-attractors approaches a new attractor given by linear inflation.
[…]
@@
As we have already seen that percent level quantum corrections can produce a notable difference in observable predictions, it would be interesting if sub-percent level corrections could produce a similar effect somewhere in this parameter space. Fig. 2 (right panel) depicts the predictions for T and E-models alongside their loop-corrected effective potentials for $\delta=0.005$. Around $\alpha\simeq 1$ we see very little departure from the predictions of the potentials without corrections, which is not surprising considering that percent level corrections were required to generate significant movement towards the central regions of the posteriors in $\xi$-attractor predictions. However, as $\alpha$ (and $r$) decrease, we begin to see far more significant departures from the uncorrected predictions. Around $\alpha\simeq 0.1$, the observables for $(n_{s},r)$ evolve rapidly towards the regions favoured by the data posteriors.
@@ III.3 Toy model with matter fields
So far, we have calculated everything at the phenomenological level, assuming radiative corrections from unspecified field content of magnitude $\delta$. Of course, the exact corrections will depend on the field content of whatever beyond the standard model physics is lurking at higher energies. To illustrate how corrections like this might arise, we consider a simple toy model of a massive spectator field. To the Lagrangian defined by Eq. (1), we add a massive field $\sigma$ with an interaction term
@@
$V_{\sigma}(\varphi,\sigma)=\frac{1}{2}\left(m_{0}^{2}+\beta f(\varphi)\right)\sigma^{2}.$ (19)
@@
where $m^{2}=V^{\prime\prime}$. For example, the effective $m^{2}_{\sigma}$ is given by $m^{2}_{\sigma}=m_{0}^{2}+\beta f(\varphi)$. Assuming that the interaction term is large $\beta\varphi^{2}\gg m^{2}_{0}$, we can write Eq. (20) and the corrected potential as,
$V_{\text{eff}}(\varphi)\simeq\lambda f^{2}(\varphi)\left[1+\frac{n\beta^{2}}{64\pi^{2}\lambda}\log\left(\frac{\varphi}{M_{\mathrm{P}}}\right)\right],$ (21)
@@
$\delta=\frac{n\beta^{2}}{64\pi^{2}\lambda}.$ (22)
Fixing $\delta=0.025$, we can then solve for $\beta$. We then explore for three different choices for all $n$ considered in Fig. 2: (i) for $\xi=1$ and $\lambda\simeq 10^{-10}$ (chosen to match the amplitude of scalar fluctuations observed in the CMB), (ii) for $\xi=10$ and $\lambda\simeq 10^{-8}$, and (iii) for $\xi=100$ and $\lambda\simeq 10^{-6}$. Furthermore, here $\chi_{\mathrm{cmb}}\simeq 6.0,5.75,5.5$ for the choices above. Both of these choices, for all $n$ considered here, result in nearly identical predictions for $(n_{s},r)$ due to their attractor behavior (i.e., for each choice of $n$ at $\xi\simeq\mathcal{O}(1)$ or greater, the predictions are insensitive to increasing $\xi$), but have different implications for the simple toy model with an additional massive field. We can then determine the effective mass $\mathcal{M}^{2}_{\sigma}\simeq\beta f(\varphi)$ in order to produce $\delta=0.025$ radiative corrections for each member of the $\xi$-attractors.
As depicted in Fig. 5, this requires very heavy masses for the $\sigma$ field in the GUT-scale range of energies, where the mass required to produce these corrections grows approximately exponentially with $n$. Interestingly, the toy model works well at smaller non-minimal coupling strengths for all $n$ and for $n=1,2$ regardless of non-minimal coupling strength. However, the model clearly begins to breakdown at larger $\xi$ and larger $n$ as the mass required to induce the considered radiative corrections in this simple toy model can approach or even exceed to the Planck mass, which indicates that it cannot be trusted in this regime as we do not have a theory of quantum gravity. This is only a toy model, but it does illustrate how one might think of the origin of the phenomenological radiative corrections considered here as well as some limitations depending on the model considered. But of course, many other models of higher energy physics are possible. See also [60] for a related discussion in the case of $\alpha$-attractors where the authors conclude that a model like this would also require GUT scale masses to generate similar radiative corrections to the ones considered in this paper.
Figure 5: The effective mass of the spectator field required to produce the radiative corrections for the $\xi$-attractors in Fig. 2 for the various $n$ values considered where the potential is given by $V(\varphi)=\lambda f^{2}(\varphi)=\lambda\varphi^{2n}$, the non-minimal coupling is $\Omega(\varphi)=1+\xi f(\varphi)=1+\xi\varphi^{n}$, and the effective mass is $\mathcal{M}^{2}_{\sigma}=\beta f(\varphi)$.
@@ IV Discussion
The recent results from ACT, if they hold up to further scrutiny, at first glance seem to be indicating that several previously favoured classes of inflationary models are notably less viable than previously thought as they lie almost entirely outside the posterior constraints on $(n_{s},r)$. However, the results here, as well as those of [28], indicate that even relatively small radiative corrections can shift $n_{s}$ into the regions indicated by the latest ACT dataset combination. This could be interpreted favourably for these models as such corrections are often seen as natural/expected given our knowledge of quantum field theory and the standard model and motivate the further exploration of radiative corrections in specific contexts with good theoretical motivations to gain insight into UV particle content. Conversely, such sensitivity to radiative corrections may imply that other models which are thought to naturally occupy a more rightward space in $n_{s}$ (an example that comes to mind is D-brane models [72]) could actually lie too far to the right if these effects are accounted for.
@@
On the other hand, a more pessimistic viewpoint is that this speaks against our ability to ever really say with definitive certainty what inflation predicts. After all, there would probably be far less interest in adding loop corrections, refining reheating predictions, investigating higher order effects, etc. if the data remained centered on the Starobinsky-Higgs values. In other words, this is yet another example where higher order effects can have a significant impact on observable predictions and inferences regarding model viability, such that one can effectively use such higher order effects to achieve almost any desired value to fit within observable constraints (another recent example includes hilltop models where it has been shown that higher order terms that stabilize the potential can likewise have a dramatic impact on what were previously considered to be definitive predictions for the models in question [73, 74, 75]). Amidst such uncertainty, what can we really say about inflation given that cosmological data seems be to unable to discriminate amongst the multitude of options regarding the nature of inflation, an uncertainty which also persists regarding the other two big open questions in cosmology (dark matter and dark energy [76, 77])?
@@
First, as has been the case for a while, it appears that non-minimal couplings (or non-trivial modifications to the kinetic sector as in the case of $\alpha$-attractors) are becoming increasingly unavoidable as they play an important role in allowing inflationary models to satisfy bounds on $r$ [78, 79, 51, 22, 80, 81, 51]. And furthermore, they are widely expected to be present at these energies. Interestingly, dark energy research concerning the current epoch of accelerating expansion seems to also be following this trend of pointing towards ever more intricate scalar field constructs [54, 53]. Second, single-field models quite generically predict a small negative running of the spectral index $\alpha_{s}\equiv dn_{s}/d\ln k$, where the value lies between $-1.8\times 10^{-3}<\alpha_{\mathrm{S}}<-9.1\times 10^{-5}$ [82]. Radiative corrections do not spoil this for any of the models considered here, so they still are within the scope of this generic prediction. The recent ACT results have also notably shifted the constraints on $\alpha_{s}$ towards the positive direction (see [24, Fig. 4]). While single field inflation models like those considered here are still compatible with the constraints, in addition to the rightward shift in $n_{s}$, this may be another hint that some of the historically more favoured models of inflation may be in tension with new data. If constraints on $\alpha_{s}$ continue to migrate towards the positive direction, this may necessitate moving away from the class of simple single-field inflation altogether. Similarly, investigations relating to generating certain features in the primordial power spectrum or in producing primordial black holes also might also have something to say regarding the ultimate viability of the more basic single-field inflation paradigm when compared with more intricate constructs within single-field modeling or multi-field models [83, 84, 85, 86, 87, 88, 89].
@@ Acknowledgements
I am very grateful to David Kaiser, Pedro Ferreira, Ioannis Gialamas, Andrei Linde, Orlando Luongo, and Antonio Delgado for illuminating discussions. I am also grateful for financial support from St. Cross College, University of Oxford.

### v3

@@ III.2 Predictions for $(n_{s},r)$
Figure 2: Comparison of $n_{s}$ and $r$ predictions for $\xi$-attractors (top left) and $\alpha$-attractor T and E-models (top right). Here we see that percent level quantum corrections can move the $\xi$-attractors off of their strong coupling attractor and comfortably within the latest ACT posteriors for the central $N=55$ number of e-foldings where we have considered values of the non-minimal coupling parameter $\log_{10}\xi\in[-8,\ 3.5]$. Additionally, we see that sub-percent level quantum corrections can induce fairly substantial shifts in the predictions for $\alpha$-attractors at small $\alpha$, where we have included the predictions spanning $N\in[50,60]$ for both T and E-models considering values of $\alpha\in[.02,4.0]$.
Figure 3: The results for $\alpha_{s}$ and $n_{s}$ (bottom) for $\xi$-attractors and $\alpha$-attractors for representative models from these classes of models ($\xi=10$ and $\alpha=0.1$) with a percent level correction spanning $N\in[50,60]$. We see $n_{s}$ shift as expected, but the radiative corrections do not induce a notable change in $\alpha_{s}$.
@@
As shown by [69, 29] for Higgs-like inflation in both the Palatini and metric variations, as $\delta$ is increased, the model begins to approach a new attractor of linear inflation $V(\varphi)\propto\varphi$. The same holds for the rest of the $\xi$-attractors considered here (see also [93, 69] for discussion of the linear attractor in this case). Choosing $\delta=0.2$ and $\log_{10}\xi\in[-8,5]$, one can see in Fig. 4 that they all begin to approach the predictions for linear inflation. As before, the higher polynomials approach the attractor more slowly in both $\delta$ and $\xi$ as their potentials are more robust against quantum corrections.
Figure 4: As the loop correction term $\delta$ is made larger, the entire family of $\xi$-attractors approaches a new attractor given by linear inflation. See also [93].
Coming now to the $\alpha$-attractor models, they behave similarly to the $\xi$-attractor models with respect to the changes in their observables when $\alpha\simeq 1$. Although not exactly equivalent due to the aforementioned difference resulting from the conformal transformation that is performed in the case of the $\xi$-attractors, percent level loop corrections can noticeably shift their predictions to be within ACT posteriors. This is not surprising as within this regime the $\alpha$-attractor potentials are very similar to the $\xi$-attractors and give almost identical predictions. However, as $\alpha\ll 1$, $r$ can be lowered arbitrarily by tuning $\alpha$ to be lower, while $n_{s}$ is independent of this choice. Thus, we will explore a wide range of potentials, choosing $\alpha\in[.02,4.0]$ which roughly spans $\log_{10}r\in[-4,-2]$ and covers most of the bounds on $r$, including going an order of magnitude below CMB-S4 projected sensitivity [105].
@@
As we have already seen that percent level quantum corrections can produce a notable difference in observable predictions, it would be interesting if $\mathcal{O}(10^{-3})$ sub-percent level corrections could produce a similar effect somewhere in this parameter space. Fig. 2 (right panel) depicts the predictions for T and E-models alongside their loop-corrected effective potentials for $\delta=0.005$. Around $\alpha\simeq 1$ we see very little departure from the predictions of the potentials without corrections, which is not surprising considering that percent level corrections were required to generate significant movement towards the central regions of the posteriors in $\xi$-attractor predictions. However, as $\alpha$ (and $r$) decrease, we begin to see far more significant departures from the uncorrected predictions. Around $\alpha\simeq 0.1$, the observables for $(n_{s},r)$ evolve rapidly towards the regions favoured by the data posteriors.
@@
While radiative corrections can clearly induce a notable shift in $n_{s}$, as also displayed in Fig. 3, introducing radiative corrections does not induce any notable changes in the values of $\alpha_{s}$ as these values remain small, negative, and right in line with the magnitudes expected from the standard Starobinsky model.
@@ III.3 Toy model with matter fields
So far, we have calculated everything at the phenomenological level, assuming radiative corrections from unspecified field content of magnitude $\delta$. Of course, the exact corrections will depend on the field content of whatever beyond the standard model physics is lurking at higher energies. To illustrate how corrections like this might arise, we consider a simple toy model of a massive scalar field coupled to the inflaton. To the Lagrangian defined by Eq. (1), we add a massive field $\sigma$ with an interaction term
@@
$V_{\sigma}(\varphi,\sigma)=\frac{1}{2}\left(m_{0}^{2}+\kappa f(\varphi)\right)\sigma^{2}.$ (20)
@@
where $m^{2}=V^{\prime\prime}$. For example, the mass is given by $m^{2}_{\sigma}=m_{0}^{2}+\kappa f(\varphi)$. Assuming that the interaction term is large $\kappa f(\varphi)\gg m^{2}_{0}$, we can write Eq. (21) and the corrected potential as,
$V_{\text{eff}}(\varphi)\simeq\lambda f^{2}(\varphi)\left[1+\frac{n\kappa^{2}}{64\pi^{2}\lambda}\log\left(\frac{\varphi}{M_{\mathrm{P}}}\right)\right],$ (22)
@@
$\delta=\frac{n\kappa^{2}}{64\pi^{2}\lambda}.$ (23)
Fixing $\delta=0.01$, we can then solve for $\kappa$. We then explore for three different choices for all $n$ considered in Fig. 2: (i) for $\xi=1$ and $\lambda\simeq 10^{-10}$ (chosen to match the amplitude of scalar fluctuations observed in the CMB) and (ii) for $\xi=10$ and $\lambda\simeq 10^{-8}$. Furthermore, here $\chi_{\mathrm{cmb}}\simeq 6.0,5.5$ for the choices above. Both of these choices, for all $n$ considered here, result in nearly identical predictions for $(n_{s},r)$ due to their attractor behavior (i.e. for each choice of $n$ at $\xi\simeq\mathcal{O}(1)$ or greater, the predictions are insensitive to increasing $\xi$), but have different implications for the simple toy model with an additional massive field. We can then determine the effective mass $\mathcal{M}^{2}_{\sigma}\simeq\kappa f(\varphi)$ in order to produce $\delta=0.01$ radiative corrections for each member of the $\xi$-attractors.
As depicted in Fig. 6, this requires very heavy masses for the $\sigma$ field around or above the GUT-scale range of energies, where the mass required to produce these corrections grows approximately exponentially with $n$. Interestingly, the toy model works well at smaller non-minimal coupling strengths for all $n$ and for $n=1,2$ regardless of non-minimal coupling strength. However, the model clearly begins to breakdown at larger $\xi$ and larger $n$ as the mass required to induce the considered radiative corrections in this simple toy model approaches the Planck mass, which indicates that it cannot be trusted in this regime as we do not have a theory of quantum gravity. This is only a toy model, but it does illustrate how one might think of the origin of the phenomenological radiative corrections considered here as well as some limitations depending on the model considered. But of course, many other models of higher energy physics are possible. See also [94] for a related discussion in the case of $\alpha$-attractors where the authors conclude that a model like this would also require GUT scale masses to generate similar radiative corrections to the ones considered in this paper.
Figure 6: The effective mass squared of the spectator field required to produce the radiative corrections for the $\xi$-attractors in Fig. 2 for the various $n$ values considered where the potential is given by $V(\varphi)=\lambda f^{2}(\varphi)=\lambda\varphi^{2n}$, the non-minimal coupling is $\Omega(\varphi)=1+\xi f(\varphi)=1+\xi\varphi^{n}$, and the effective mass is $\mathcal{M}_{\sigma}=\sqrt{\kappa f(\varphi)}$.
@@ IV Discussion
The recent results from ACT, if they hold up to further scrutiny, at first glance seem to be indicating that several previously favoured classes of inflationary models are notably less viable than previously thought as they lie almost entirely outside the posterior constraints on $(n_{s},r)$. However, the results here, as well as those of [29], indicate that even relatively small radiative corrections can shift $n_{s}$ into the regions indicated by the latest ACT dataset combination. This could be interpreted favourably for these models as such corrections are seen by some as natural/expected given our knowledge of quantum field theory/standard model particle physics (assuming that the UV completion is one that would lead to non-negligible radiative corrections) and motivate the further exploration of radiative corrections in specific contexts with good theoretical motivations to gain insight into UV particle content. In particular, it may motivate finding or exploring particular UV completions that would imply corrections of these magnitudes.
@@
On the other hand, if UV physics is such that we should expect there to be non-negligible radiative corrections to inflation, this would imply a pessimism regarding our ability to ever really say with definitive certainty what inflation predicts. After all, there would probably be far less interest in adding loop corrections, refining reheating predictions, investigating higher order effects, etc. if the data remained centered on the Starobinsky values. In other words, this is yet another example where higher order effects can have a non-negligible impact on observable predictions and inferences regarding model viability, such that one can effectively use such higher order effects to achieve almost any desired value to fit within observable constraints (another recent example includes hilltop models where it has been shown that higher order terms that stabilize the potential can likewise have a dramatic impact on what were previously considered to be definitive predictions for the models in question [108, 109, 110]). Amidst such uncertainty, what can we really say about inflation given that cosmological data seems be to unable to discriminate amongst the multitude of options regarding the nature of inflation, an uncertainty which also persists regarding the other two big open questions in cosmology (dark matter and dark energy [111, 112])?
First, as has been the case for a while, it appears that non-minimal couplings (or non-trivial modifications to the kinetic sector as in the case of $\alpha$-attractors) are becoming increasingly unavoidable as they play an important role in allowing inflationary models to satisfy bounds on $r$ [113, 114, 78, 21, 115, 116, 78, 117]. And furthermore, they are widely expected to be present at these energies. Interestingly, dark energy research concerning the current epoch of accelerating expansion seems to also be following this trend of pointing towards ever more intricate scalar field constructs [81, 80]. So, all things considered, observations seem to be consistently pushing us towards more complicated sections of the single-field model space.
@@
Second, there are some signals that should shed light on the viability of the single-field paradigm. For example, single-field models quite generically predict a small negative running of the spectral index $\alpha_{s}\equiv dn_{s}/d\ln k$, where the value lies between $8\times 10^{-3}<\alpha_{\mathrm{S}}<-9.1\times 10^{-5} [11$8]. Radiative corrections do not spoil this for any of the models considered here, so they still are within the scope of this generic prediction. The recent ACT results have also notably shifted the constraints on $\alpha_{s}$ towards the positive direction (see [25, Fig. 4]). While single field inflation models like those considered here are still compatible with the constraints, in addition to the rightward shift in $n_{s}$, this may be another hint that some of the historically more favoured models of inflation may be in tension with new data. If constraints on $\alpha_{s}$ continue to migrate towards the positive direction, this may necessitate moving away from the class of simple single-field inflation altogether. Similarly, investigations relating to generating particular features in the primordial power spectrum, producing primordial black holes, and certain signals in the stochastic gravitational wave background also might also have something to say regarding the ultimate viability of the more basic single-field inflation paradigm when compared with more intricate constructs within single-field modeling, multi-field models, or even more exotic alternatives [119, 120, 121, 122, 123, 124, 125, 126].
@@ Acknowledgements
I am very grateful to David Kaiser, Pedro Ferreira, Ioannis Gialamas, Andrei Linde, Orlando Luongo, Sunny Vagnozzi, and Antonio Delgado for illuminating discussions. I am also grateful for financial support from St. Cross College, University of Oxford.